# Bayesian spatio-temporal weighted regression for integrating missing and misaligned environmental data

### v1

@@ Abstract
Estimating environmental exposures from multi-source data is central to public health research and policy. Integrating data from satellite products and ground monitors are increasingly used to produce exposure surfaces. However, spatio-temporal misalignment often induced from missing data introduces substantial uncertainty and reduces predictive accuracy. We propose a Bayesian weighted predictor regression framework that models spatio-temporal relationships when predictors are observed on irregular supports or have substantial missing data, and are not concurrent with the outcome. The key feature of our model is a spatio-temporal kernel that aggregates the predictor over local space-time neighborhoods, built directly into the likelihood, eliminating any separate gap-filling or forced data alignment stage. We introduce a numerical approximation using a Voronoi-based spatial quadrature combined with irregular temporal increments for estimation under data missingness and misalignment. We showed that misspecification of the spatial and temporal lags induced bias in the mean and parameter estimates, indicating the need for principled parameter selection. Simulation studies confirmed these theoretical findings, where careful tuning was critical to control bias and achieve accurate prediction, while the proposed quadrature performed well under severe missingness. As an illustrative application, we estimated fine particulate matter (PM2.5) in northern California using satellite-derived aerosol optical depth (AOD) and wildfire smoke plume indicators. Relative to a traditional collocated linear model, our approach improved out-of-sample predictive performance (over 50% increase in R2), reduced uncertainty, and yielded robust temporal predictions and spatial surface estimation. Our framework is extensible to additional spatio-temporally varying covariates and other kernel families.
Keywords Spatio-temporal regression $\cdot$ Kernel-averaged predictor $\cdot$ Missing and misaligned data $\cdot$ PM2.5 mapping
@@ 1 Introduction
Environmental exposures derived from multi-source data play a central role in public health. Accurate exposure estimates are needed to quantify population-level risks, guide interventions, and inform regulatory action [Clark et al., 2025]. Data streams such as satellite observations, ground monitors, and physical model outputs provide complementary information, but they often differ in spatial resolution, temporal frequency, and coverage [Lolli, 2025]. Integrating these sources is essential for many applications but can be methodologically challenging [Forlani et al., 2020]. Missing observations, spatial or temporal misalignment, and nonstationary associations between predictors and outcomes introduce uncertainty, limiting predictive reliability and the robustness of scientific inference.
@@
Air quality research provides a clear example of these challenges. Exposure to fine particulate matter (PM2.5) is a leading risk factor for respiratory and cardiovascular disease and premature mortality worldwide, contributing to an estimated 8.1 million deaths globally in 2021 [Health Effects Institute, 2024]. While ground-based monitoring networks provide reliable measurements, they are spatially sparse and unevenly distributed [Zhang et al., 2021], limiting their ability to capture fine-scale spatial variability. Satellite-derived aerosol optical depth (AOD), available from instruments like the Moderate resolution Imaging Spectroradiometer (MODIS), offer global spatial coverage and have been widely used as proxies for surface PM2.5 [Diao et al., 2019]. However, AOD retrievals are often missing due to cloud cover, snow (bright surfaces), or algorithmic failures, and when available, they are not always co-located or spatially aligned with ground monitors, leading to incomplete or misaligned data.
@@
To address these limitations, statistical approaches have been developed to model the PM2.5–AOD relationship and integrate additional sources of data such as meteorology and land use characteristics. Early methods relied on linear regression [Wang and Christopher, 2003, Chu et al., 2003] and were later extended through multiple linear regression and generalized additive models to incorporate meteorological covariates [van Donkelaar et al., 2006, Liu et al., 2009, Franklin et al., 2017]. In these studies, spatial misalignment was commonly addressed by interpolating or averaging AOD onto regular grids first, then modeling the PM-AOD relationship. Machine learning and deep learning have come to the forefront in interpolation efforts as well, but explicit spatio-temporal trends are not generally incorporated in these approaches [Li et al., 2020]. Multi-stage spatio-temporal models have been proposed to estimate daily PM2.5 using neighbouring observations where AOD is missing, [Pu and Yoo, 2020]; however, these approaches remain limited in their ability to capture nonstationarity and locally varying relationships.
@@
Local regression techniques provide one pathway to address this gap. Geographically weighted regression (GWR) [Brunsdon et al., 1998] and its spatio-temporal extension, GTWR [Fotheringham et al., 2015] apply location- and time-specific weights, allowing regression coefficients to vary across both space and time. GTWR has been shown to improve predictive performance for PM2.5 mapping by capturing nonstationary AOD–PM2.5 relationships [He and Huang, 2018, Hu et al., 2013], producing fine-resolution exposure surfaces that reveal gradients and hotspots. Yet, these models still rely on concurrent predictor–outcome pairs, making them vulnerable to the systematic missingness and misalignment common in environmental datasets.
@@
Other methodological advances in spatial modeling have aimed to extend beyond strictly concurrent data. Kernel-averaged predictor frameworks [Heaton and Gelfand, 2011, Heaton and Gelfand, 2012] smoothed predictors over a spatial or spatio-temporal buffer, allowing nearby information to inform estimates even when direct matches are unavailable. While promising, these methods can be sensitive to the choice of kernel size and lag parameters, and their robustness to missing data remains unclear. This is a pressing concern in many applications, where monitoring networks are sparse, satellite retrievals are often intermittent, and episodic events such as wildfire smoke plumes create rapidly evolving exposure patterns.
@@
In this paper, we build on the developments of [Heaton and Gelfand, 2012] by proposing a new method of estimation for the weighted predictor regression framework, tailored to settings with missing or misaligned data. Our Bayesian formulation allows flexible spatial and temporal lag structures, enabling regression even when predictors are not observed concurrently with the outcome. The framework accommodates varying buffer sizes and temporal lag specifications, and we evaluate its robustness to model misspecification as well as its performance under substantial missingness in the predictors. We applied the approach to estimate PM2.5 in Northern California from satellite AOD and wildfire smoke plumes, quantifying local associations and demonstrating gains in predictive performance relative to traditional concurrent data models. The paper is organized as follows: Section 2 describes the methodology, provides theoretical bounds for bias induced by model misspecification, and outlines model fitting procedures. Section 3 presents a simulation study evaluating model sensitivity and robustness. Section 4 applies the methods to study spatio-temporal dependencies between daily PM2.5 and AOD during a year where there was strong influence of wildfires on air quality in the region.
@@ 2 Method
We denote a spatio-temporal process by $\{Y(\mathbf{s},t):\mathbf{s}\in\mathbb{R}^{d},t\in\mathbb{R}\}$ where $Y(\mathbf{s},t)$ represents an observation at location $\mathbf{s}$ and time $t$ and assume that $Y(\mathbf{s},t)$ depends on another spatio-temporal temporal process $\{X(\mathbf{u},\tau):\mathbf{u}\in\mathbb{R}^{d},\tau\in\mathbb{R}\}$. In a general setting, the proposed model, similar to [Heaton and Gelfand, 2012], is given by
@@
This framework captures spatio-temporal variation by accounting for the local weighting of nearby predictors and recent time points, as well as the regression effect. These two sets of parameters play different roles: the (kernel) decay parameters $\xi(\mathbf{s},t)$ determine the spatial extent over which the predictors can influence the parameters, whereas the regression parameters $\beta_{j}(\mathbf{s},t)$ govern the strength of the association between the weighted predictor and the response. We model both sources of non-stationarity using smooth functions or covariates observed at the corresponding locations and times:
$\beta_{j}(\mathbf{s},t)=\sum b_{jk}f_{k}(s,t)\textrm{ and }\mathbb{E}(\xi(\mathbf{s},t))=\sum_{g=1}^{G}e_{g}v_{g}(s,t),\;\xi(\mathbf{s},t)\sim\mathcal{G}(\cdot)$
@@ 2.1 Choice of radius and time lag
The choice of $r$ and $q$ must be treated as a model selection problem in which values that maximize goodness-of-fit measures such as R2 and/or minimize Akaike Information Criteria (AIC) and mean-squared error (MSE) are selected. [Heaton and Gelfand, 2011] choose $r$ based on the effective spatial range of the predictor. However, the effect of misspecifying $r$ and $q$ on parameter estimation and predictive performance remains unexplored. In this work, we focus on kernel functions of the form
@@
Proposition 1 implies that the discrepancy between the true and misspecified means is influenced not only by the behaviour of the kernel functions but also by the magnitude of the process outside the true spatial–temporal window.
@@
From collorary 1, if $\delta_{1},\delta_{2}\to 0$, then the bound only depends on the behaviour of $X(\mathbf{u},t-l)$ i.e. as long as $X(\mathbf{u},t-l)$ does not contain extremes, the error bound $\to 0$. In this setting, one might consider selecting $\tilde{r}$ and $\tilde{q}$ such that $\mathcal{B}(\tilde{r})\times[0,\tilde{q}]=\mathcal{D}\times\mathcal{T}$, thereby covering the entire spatio-temporal domain under consideration. While this choice appears reasonable in theory, it is often computationally prohibitive in practice. For example, on a regular 1 km² daily grid, the number of data points required to cover a region of radius $\mathcal{B}(\tilde{r})\times[0,\tilde{q}$] is $\pi\tilde{r}^{2}\tilde{q}$, which grows quadratically with $\tilde{r}$. Thus, it is recommended that model selection be based on performance assessed under a set of varying spatial and temporal lag values.
@@ 2.2.1 Numerical approximation of spatio-temporal integral
When data are missing or misaligned, this estimation is more challenging. [Heaton and Gelfand, 2011] use Monte Carlo integration and [Heaton and Gelfand, 2012] project $X(\mathbf{s},t)$ onto a regular grid and approximate the integral by summing pointwise products of the discretised covariate with its weights after fitting a latent Gaussian process - an approach that can be computationally expensive for finely spaced data due to covariance estimation. We instead use a numerical approximation for the weighted predictor function:;
@@
A Voronoi tessellation is a computational technique that partitions a metric space based on proximity to a given set of points [Okabe et al., 2009]. At time $i$, given a finite set of distinct points given by $\mathbf{z}_{ni}$ which represents the locations at which we have observed data within the buffer of location $n$, the Voronoi cell $V_{ni,k}$ associated with location $z_{ni,k}$ is defined as:
@@
Because $Y$ and $X$ are typically collected on different space-time supports, the choice of $r$ and $q$ directly affects $M_{ni}$, the number of covariate values contributing to $W(\mathbf{s}_{n},t_{i})$. In applications such as PM2.5 estimation from satellite AOD, there can be substantial missing data in $X(\mathbf{v},\tau)$ that yield $M_{ni}=0$ (or very small), in which case $Y(\mathbf{s}_{n},t_{i})$ is poorly approximated. In this case, $Y(\mathbf{s}_{n},t_{i})$ should be removed to avoid large approximation error. In Appendix A Figure 7, we illustrate the Voronoi tessellation of the buffer on two different days. The tessellation depends on the location of the observed AOD data.
@@ 2.2.2 Bayesian inference
A Metropolis-within-Gibbs sampler is used to perform inference. We assume a Gaussian for prior $\mathbf{\beta}\sim\mathcal{N}(0,\sigma^{2}_{b})$. This choice of prior is conditionally conjugate under a Gaussian process assumption for $\varepsilon(\mathbf{s},t)\sim\mathcal{GP}(0,\sigma^{2}C(\mathbf{s},\mathbf{s}^{\prime},t,t^{\prime}))$. A conditionally conjugate prior $\mathcal{IG}(a_{b},b_{b})$ is assumed for $\sigma_{b}^{2}$. Similarly, a conditionally conjugate non-informative inverse-Gamma prior is assumed for $\sigma^{2}\sim\mathcal{IG}(a,b)$. The prior for $\mathbf{\xi}(\mathbf{s},t)$ depends on the choice of covariance function and assumed model. In general for $\mathbf{\xi}(\mathbf{s},t)=\xi$ , we propose an informative Gamma $\sim\mathcal{G}(c_{1},c_{2})$ for non-spatio-temporal scale parameters governing the kernel functions particularly in the presence of missing data. A natural choice for $c_{1}$ and $c_{2}$ is based on the expected rate of decay over which the predictor is allowed to extend. Under this prior specification, a Metropolis-Hasting step is used to sample from the posterior distribution. Alternative choice of priors for different parameterizations of the scale parameters is discussed further in Section 4. Full details of the Metropolis-within-Gibbs sampler implementation are available in Appendix D.
@@ 3 Simulation Study
$\mu_{t}(\mathbf{s})=$ $\exp\left(-\left(\frac{(s_{1}-s_{1,1,t})^{2}}{2\cdot 50^{2}}+\frac{(s_{2}-s_{2,1,t})^{2}}{2\cdot 10^{2}}\right)\right)+1.5\exp\left(-\left(\frac{(s_{1}-s_{1,2,t})^{2}}{2\cdot 70^{2}}+\frac{(s_{2}-s_{2,2,t})^{2}}{2\cdot 15^{2}}\right)\right)$
$+0.5\exp\left(-\left(\frac{(s_{1}-s_{1,3,t})^{2}}{2\cdot 40^{2}}+\frac{(s_{2}-s_{2,3,t})^{2}}{2\cdot 8^{2}}\right)\right)$
@@
where $\varepsilon\sim\mathcal{N}(0,0.25^{2})$ and $c_{1}$ and $c_{2}$ are normalizing constants for the spatial and temporal kernels so that each integrates to 1 over its support. In this model specification, only predictors within a 10 km radius and the most recent 5 time lags affect the response (r=10,q=5r=10,\;q=5). Illustrations of the simulated surface and predictor process are provided in Appendix B. As shown in Figure 9, the spatio-temporal weights rapidly decay ($\to 0$) for $r>10$ and $l>4$.
We generated 100 replicated data sets on N = 100 locations and T = 153 time points. We retained N = 25 locations for model fitting and the remaining out-of-sample to compare predictive performance. This choice reflects the sparseness of ground-level stations in practice. Using the fully observed predictor process, we fit (i) a model with the true lag values $(r,q)=(10,5)$, (ii) a truncated misspecified model with $\tilde{r}=5<r$ and $\tilde{q}=2<q$ (iii) an extended misspecified model with $\tilde{r}=15>r$ and $\tilde{q}=8>q$. The covariate process $X_{t}(\mathbf{s})$ is well-behaved (no heavy tails or extremes). To assess the approximation under irregular sampling, we also simulated datasets with missing values that mimic the application’s missingness pattern observed (described below).For inference, we found that using a non-informative inverse-Gamma prior, $\sigma^{2}\sim\mathcal{IG}(a_{1},b_{1}),$ does not pose issues for parameter estimation or identifiability when the data are complete. However, in the presence of missing data, employing an informative prior helps improve both identifiability and sampler mixing.
@@
Figure  1 illustrates the effect of model misspecification on parameter estimation. When $\tilde{r}<r$ and $\tilde{q}<q$, the model tends to overestimate parameters and exhibit greater variability than the correctly specified model. In particular, we observe large biases in the estimated scale parameter of the spatial kernel function. This pattern appeared both with full data setting and under missingness. A larger value or $\xi_{1}$ corresponds to a more uniform weighting across the radius (i.e. slower decay). Consequently, selecting a smaller radius can lead the model to overestimate the influence of predictors within $\tilde{r}$ to compensate for omitted contributions from $\mathcal{A}_{\tilde{r},r}(\mathbf{s})$. Similarly, the effect of parameter $\beta_{1}$ is overestimated to offset the mean bias induced by the reduced integration domain.
With missing data, the numerical approximation introduces some bias; even with the true $(r,q)$, parameter estimates are positively biased and the bias is typically exacerbated under model misspecification. The top left panel in Figure 2 displays the distribution of the noise parameter $\sigma^{2}$. Misspecification of the mean process results in systematically higher estimates of $\sigma^{2}$. In the presence of missing data, estimated $\sigma^{2}$ is generally larger, reflecting increased uncertainty.
Figure 2: Estimation of $\sigma^{2}$ and out-of-sample performance of simulated data measures using the $R^{2}$ metric, mean-squared error and coverage of the 95% prediction interval. Red boxplots indicate the results obtained using true radius and time lag values $(r=10,q=5)$, while blue $(\tilde{r}=15,\tilde{q}=8)$ and green $(\tilde{r}=5,\tilde{q}=2)$ illustrate the effect of model misspecification on predicitive performance.
Despite these biases, the $R^{2}$ values remain relatively high, particularly when a reduced region of integration is selected. This pattern persists in the presence of missing data. Similarly, the mean-squared error is low relative to the performance under the true parameters, although misspecified models exhibit greater variability in out-of-sample performance. Under the full data setting, coverage of the reduced model is lower, with prediction interval widths ranging from 2.81 to 6.30, compared to the wider intervals of the extended model (2.75 to 8.44). Finally, even with missing data, out-of-sample performance remains strong, with coverage exceeding 95%, reflecting wider prediction intervals and higher uncertainty estimates.

### v2

# Methodology, Formal Analysis, Software, Validation, Visualization, Writing – original draft

@@ Abstract
Estimating environmental exposures from multi-source data is central to public health research and policy. Integrating data from satellite products and ground monitors are increasingly used to produce exposure surfaces. However, spatio-temporal misalignment often induced from missing data introduces substantial uncertainty and reduces predictive accuracy. We propose a Bayesian weighted predictor regression framework that models spatio-temporal relationships when predictors are observed on irregular supports or have substantial missing data, and are not concurrent with the outcome. The key feature of our model is a spatio-temporal kernel that aggregates the predictor over local space-time neighborhoods, built directly into the likelihood, eliminating any separate gap-filling or forced data alignment stage. We introduce a numerical approximation using a Voronoi-based spatial quadrature combined with irregular temporal increments for estimation under data missingness and misalignment. We showed that misspecification of the spatial and temporal lags induced bias in the mean and parameter estimates, indicating the need for principled parameter selection. Simulation studies confirmed these findings, where careful tuning was critical to control bias and achieve accurate prediction, while the proposed quadrature performed well under severe missingness. As an illustrative application, we estimated fine particulate matter (PM2.5) in northern California using satellite-derived aerosol optical depth (AOD) and wildfire smoke plume indicators. Relative to a traditional collocated linear model, our approach improved out-of-sample predictive performance, reduced uncertainty, and yielded robust temporal predictions and spatial surface estimation. Our framework is extensible to additional spatio-temporally varying covariates and other kernel families.
keywords:
[…]
@@ 1 Introduction
Environmental exposures derived from multi-source data play a central role in public health. Accurate exposure estimates are needed to quantify population-level risks, guide interventions, and inform regulatory action (clark2025). Data streams such as satellite observations, ground monitors, and physical model outputs provide complementary information, but they often differ in spatial resolution, temporal frequency, and coverage (lolli2025). Integrating these sources is essential for many applications but can be methodologically challenging (forlani2020). Missing observations, spatial or temporal misalignment, and nonstationary associations between predictors and outcomes introduce uncertainty, limiting predictive reliability and the robustness of scientific inference.
@@
Air quality research provides a clear example of these challenges. Exposure to fine particulate matter (PM2.5) is a leading risk factor for respiratory and cardiovascular disease and premature mortality worldwide, contributing to an estimated 8.1 million deaths globally in 2021 (health_effects_institute_state_2024). While ground-based monitoring networks provide reliable measurements, they are spatially sparse and unevenly distributed (zhang_satellite_2021), limiting their ability to capture fine-scale spatial variability. Satellite-derived aerosol optical depth (AOD), available from instruments like the Moderate resolution Imaging Spectroradiometer (MODIS), offer global spatial coverage and have been widely used as proxies for surface PM2.5 (diao2019). However, AOD retrievals are often missing due to cloud cover, snow (bright surfaces), or algorithmic failures, and when available, they are not always co-located or spatially aligned with ground monitors, leading to incomplete or misaligned data.
@@
To address these limitations, statistical approaches have been developed to model the PM2.5–AOD relationship and integrate additional sources of data such as meteorology and land use characteristics. Early methods relied on linear regression (wang_intercomparison_2003; chu_global_2003) and were later extended through multiple linear regression and generalized additive models to incorporate meteorological covariates (van_donkelaar_estimating_2006; liu_estimating_2009; franklin2017). In these studies, spatial misalignment was commonly addressed by interpolating or averaging AOD onto regular grids first, then modeling the PM-AOD relationship. Machine learning and deep learning have come to the forefront in interpolation efforts as well, but explicit spatio-temporal trends are not generally incorporated in these approaches (li2020). Multi-stage spatio-temporal models have been proposed to estimate daily PM2.5 using neighbouring observations where AOD is missing, (pu_spatio-temporal_2020); however, these approaches remain limited in their ability to capture nonstationarity and locally varying relationships.
@@
Local regression techniques provide one pathway to address this gap. Geographically weighted regression (GWR) (brunsdon_geographically_1998) and its spatio-temporal extension, GTWR (fotheringham_geographical_2015) apply location- and time-specific weights, allowing regression coefficients to vary across both space and time. GTWR has been shown to improve predictive performance for PM2.5 mapping by capturing nonstationary AOD–PM2.5 relationships (he_satellite-based_2018; hu_estimating_2013), producing fine-resolution exposure surfaces that reveal gradients and hotspots. Yet, these models still rely on concurrent predictor–outcome pairs, making them vulnerable to the systematic missingness and misalignment common in environmental datasets.
@@
Other methodological advances in spatial modeling have aimed to extend beyond strictly concurrent data. Kernel-averaged predictor frameworks (heaton_spatial_2011; heaton_kernel_2012) smoothed predictors over a spatial or spatio-temporal buffer, allowing nearby information to inform estimates even when direct matches are unavailable. While promising, these methods can be sensitive to the choice of kernel size and lag parameters, and their robustness to missing data remains unclear. This is a pressing concern in many applications, where monitoring networks are sparse, satellite retrievals are often intermittent, and episodic events such as wildfire smoke plumes create rapidly evolving exposure patterns.
@@
In this paper, we build on the developments of heaton_kernel_2012 by proposing a new method of estimation for the weighted predictor regression framework, tailored to settings with missing or misaligned data. Our Bayesian formulation allows flexible spatial and temporal lag structures, enabling regression even when predictors are not observed concurrently with the outcome. The framework accommodates varying buffer sizes and temporal lag specifications, and we evaluate its robustness to model misspecification as well as its performance under substantial missingness in the predictors. We applied the approach to estimate PM2.5 in Northern California from satellite AOD and wildfire smoke plumes, quantifying local associations and demonstrating gains in predictive performance relative to traditional concurrent data models. The paper is organized as follows: Section 2 describes the methodology, provides theoretical bounds for bias induced by model misspecification, and outlines model fitting procedures. Section 3 presents a simulation study evaluating model sensitivity and robustness. Section 4 applies the methods to study spatio-temporal dependencies between daily PM2.5 and AOD during a year where there was strong influence of wildfires on air quality in the region.
@@ 2 Method
We denote a spatio-temporal process by $\{Y(\mathbf{s},t):\mathbf{s}\in\mathbb{R}^{d},t\in\mathbb{R}\}$ where $Y(\mathbf{s},t)$ represents an observation at location $\mathbf{s}$ and time $t$ and assume that $Y(\mathbf{s},t)$ depends on another spatio-temporal temporal process $\{X(\mathbf{u},\tau):\mathbf{u}\in\mathbb{R}^{d},\tau\in\mathbb{R}\}$. In a general setting, the proposed model, similar to heaton_kernel_2012, is given by
@@
This framework captures spatio-temporal variation by accounting for the local weighting of nearby predictors and recent time points, as well as the regression effect. These two sets of parameters play different roles: the (kernel) decay parameters $\xi(\mathbf{s},t)$ determine the spatial extent over which the predictors can influence the parameters, whereas the regression parameters $\beta_{1}(\mathbf{s},t)$ govern the strength of the association between the weighted predictor and the response. We model both sources of non-stationarity using smooth functions or covariates observed at the corresponding locations and times:
$\beta_{1}(\mathbf{s},t)=\sum b_{k}f_{k}(s,t)\textrm{ and }\mathbb{E}(\xi(\mathbf{s},t))=\sum_{g=1}^{G}e_{g}v_{g}(s,t),\;\xi(\mathbf{s},t)\sim\mathcal{F}(\cdot)$
@@ 2.1 Choice of radius and time lag
The choice of $r$ and $q$ must be treated as a model selection problem in which values that maximize goodness-of-fit measures such as R2 and/or minimize Akaike Information Criteria (AIC) and mean-squared error (MSE) are selected. heaton_spatial_2011 choose $r$ based on the effective spatial range of the predictor. However, the effect of misspecifying $r$ and $q$ on parameter estimation and predictive performance remains unexplored. In this work, we focus on kernel functions of the form
@@
Proposition 1 implies that the discrepancy between the true and misspecified means is influenced not only by the behaviour of the kernel functions but also by the magnitude of the process outside the true spatial–temporal window. The proof of this proposition can be found under Appendix C of the Supplementary Material.
@@
From collorary 1, if $\delta_{1},\delta_{2}\to 0$, then the bound only depends on the behaviour of $X(\mathbf{u},t-l)$ i.e. as long as $X(\mathbf{u},t-l)$ does not contain extremes, the error bound $\to 0$. In this setting, one might consider selecting $\tilde{r}$ and $\tilde{q}$ such that $\mathcal{B}(\tilde{r})\times[0,\tilde{q}]=\mathcal{D}\times\mathcal{T}$, thereby covering the entire spatio-temporal domain under consideration. While this choice appears reasonable in theory, it is often computationally prohibitive in practice. For example, on a regular 1 km² daily grid, the number of data points required to cover a region of radius $\mathcal{B}(\tilde{r})\times[0,\tilde{q}$] is $\pi\tilde{r}^{2}\tilde{q}$, which grows quadratically with $\tilde{r}$. Thus, it is recommended that model selection be based on out-of-sample performance assessed under a set of varying spatial and temporal lag values.
@@ 2.2.1 Numerical approximation of spatio-temporal integral
When data are missing or misaligned, this estimation is more challenging. heaton_spatial_2011 use Monte Carlo integration and heaton_kernel_2012 project $X(\mathbf{s},t)$ onto a regular grid and approximate the integral by summing pointwise products of the discretised covariate with its weights after fitting a latent Gaussian process - an approach that can be computationally expensive for finely spaced data due to covariance estimation. We instead use a numerical approximation for the weighted predictor function:;
@@
A Voronoi tessellation is a computational technique that partitions a metric space based on proximity to a given set of points okabe_spatial_2009. At time $i$, given a finite set of distinct points given by $\mathbf{z}_{ni}$ which represents the locations at which we have observed data within the buffer of location $n$, the Voronoi cell $V_{ni,k}$ associated with location $z_{ni,k}$ is defined as:
@@
Because $Y$ and $X$ are typically collected on different space-time supports, the choice of $r$ and $q$ directly affects $M_{ni}$, the number of covariate values contributing to $W(\mathbf{s}_{n},t_{i})$. In applications such as PM2.5 estimation from satellite AOD, there can be substantial missing data in $X(\mathbf{v},\tau)$ that yield $M_{ni}=0$ (or very small), in which case $Y(\mathbf{s}_{n},t_{i})$ is poorly approximated. In this case, $Y(\mathbf{s}_{n},t_{i})$ should be removed to avoid large approximation error. In Appendix A, we illustrate the Voronoi tessellation of the buffer on two different days. The tessellation depends on the location of the observed AOD data. An advantage of this method is that it accounts for regions where AOD is unobserved by approximating them using the nearest observed values, rather than treating them as missing, as is often assumed in traditional collocation approaches.
@@ 2.2.2 Bayesian inference
A Metropolis-within-Gibbs sampler is used to perform inference. We assume a Gaussian for prior $b_{k}\sim\mathcal{N}(0,\sigma^{2}_{b})$. This choice of prior is conditionally conjugate under a Gaussian process assumption for $\varepsilon(\mathbf{s},t)\sim\mathcal{GP}(0,\sigma^{2}C(\mathbf{s},\mathbf{s}^{\prime},t,t^{\prime}))$. A conditionally conjugate prior $\mathcal{IG}(a_{b},b_{b})$ is assumed for $\sigma_{b}^{2}$. Similarly, a conditionally conjugate non-informative inverse-Gamma prior is assumed for $\sigma^{2}\sim\mathcal{IG}(a,b)$. The prior for $\mathbf{\xi}(\mathbf{s},t)$ depends on the choice of covariance function and assumed model. In general for $\mathbf{\xi}(\mathbf{s},t)=\xi$ , we propose an informative Gamma $\sim\mathcal{G}(c_{1},c_{2})$ for non-spatio-temporal scale parameters governing the kernel functions particularly in the presence of missing data. A natural choice for $c_{1}$ and $c_{2}$ is based on the expected rate of decay over which the predictor is allowed to extend. Under this prior specification, a Metropolis-Hasting step is used to sample from the posterior distribution. Alternative choice of priors for different parameterizations of the scale parameters is discussed further in Section 4. Full details of the Metropolis-within-Gibbs sampler implementation are available in Appendix D.
@@ 3 Simulation Study
$\mu_{t}(\mathbf{s})=$ $\exp\left(-\left(\frac{(s_{1}-s_{1,1,t})^{2}}{2\cdot 50^{2}}+\frac{(s_{2}-s_{2,1,t})^{2}}{2\cdot 10^{2}}\right)\right)+1.5\exp\left(-\left(\frac{(s_{1}-s_{1,2,t})^{2}}{2\cdot 70^{2}}+\frac{(s_{2}-s_{2,2,t})^{2}}{2\cdot 15^{2}}\right)\right)+$ (4)
$0.5\exp\left(-\left(\frac{(s_{1}-s_{1,3,t})^{2}}{2\cdot 40^{2}}+\frac{(s_{2}-s_{2,3,t})^{2}}{2\cdot 8^{2}}\right)\right)$ (5)
@@
where $\varepsilon\sim\mathcal{N}(0,0.25^{2})$ and $c_{1}$ and $c_{2}$ are normalizing constants for the spatial and temporal kernels so that each integrates to 1 over its support. In this model specification, only predictors within a 10 km radius and the most recent 5 time lags affect the response (r=10,q=5r=10,\;q=5). As shown in the illustrations of the simulated surface and predictor process provided in Appendix B, the spatio-temporal weights rapidly decay ($\to 0$) for $r>10$ and $q>3$. Finally, we set $\beta_{0}=10$ and $\beta_{1}=0.5$.
We generated 100 replicated data sets on N = 100 locations and T = 153 time points. We retained N = 25 locations for model fitting and the remaining out-of-sample to compare predictive performance. This choice reflects the sparseness of ground-level stations in practice. Using the fully observed predictor process, we fit (i) a model with the true lag values $(r,q)=(10,5)$, (ii) a truncated misspecified model with $\tilde{r}=5<r$ and $\tilde{q}=2<q$ (iii) an extended misspecified model with $\tilde{r}=15>r$ and $\tilde{q}=8>q$. The covariate process $X_{t}(\mathbf{s})$ is well-behaved (no extremes). To assess the approximation under irregular sampling, we also simulated datasets with missing values that mimic the application’s missingness pattern observed (described in Figure 3). For inference, we found that using a non-informative inverse-Gamma prior, $\sigma^{2}\sim\mathcal{IG}(a_{1},b_{1}),$ does not pose issues for parameter estimation or identifiability when the data are complete. However, in the presence of missing data, employing an informative prior helps improve both identifiability and sampler mixing.
@@
Figure  1 illustrates the effect of model misspecification on parameter estimation. When $\tilde{r}<r$ and $\tilde{q}<q$, the model is more prone to overestimating parameters. In particular, we observe large biases in the estimated scale parameter of the spatial kernel function. This pattern appeared both with full data setting and under missingness. A larger value or $\xi_{1}$ corresponds to a more uniform weighting across the radius (i.e. slower decay). Consequently, selecting a smaller radius can lead the model to overestimate the influence of predictors within $\tilde{r}$ to compensate for omitted contributions from $\mathcal{A}_{\tilde{r},r}(\mathbf{s})$. Similarly, the effect of parameters $\beta_{0}$ and $\beta_{1}$ is overestimated to offset the mean bias induced by the reduced integration domain. However, $\xi_{2}$ is estimated more consistently because the true data-generating process assumes a faster temporal decay than the lag parameter used to generate the data. However, because estimates of the scale parameters are sporadic, stronger priors should be favoured if model inference and interpretability are of primary concern. With missing data, the numerical approximation introduces some bias; even with the true $(r,q)$, parameter estimates are positively biased and the bias is typically exacerbated under model misspecification.
Figure 2: Out-of-sample performance of simulated data measured using the mean deviation of the predictions from the true values, and coverage of the 95% prediction interval. Red boxplots indicate the results obtained using true radius and time lag values $(r=10,q=5)$, while blue $(\tilde{r}=15,\tilde{q}=8)$ and green $(\tilde{r}=5,\tilde{q}=2)$ illustrate the effect of model misspecification on predictive performance.
Under the full data setting, the out-of-sample deviations ($\hat{Y}(\mathbf{s},t)-Y(\mathbf{s},t)$) results in Figure 2 align with Corollary 1. When the kernel functions decay to zero within the true region of integration, selecting larger $\tilde{r}$ and $\tilde{q}$ reduces the mean bias compared to a smaller integration region. Additionally, predictions are underestimated when choosing smaller lags with lower coverage of the 95% prediction intervals. In the presence of missing data, we observe similar patterns in the deviations of the predictions from the true value. Most importantly, our proposed numerical approximation produces robust model predictions with high coverage of the prediction intervals. These results show that in practice choosing sufficiently large lag values ensures both reliable parameter estimation and predictions.